OpenStudy (anonymous):

Solve Exactly: log2(3x+20) - log2 =3

OpenStudy (anonymous):

thank you

OpenStudy (tkhunny):

Are you sure we have a proper problem statement?

OpenStudy (anonymous):

I believe so. It's from a practice quiz my teacher gave me. But I couldn't find any similar problems like it in my book.

OpenStudy (tyteen4a03):

@dinosgorawr Please do not give out answers without explanations. It does not help the user learn, and is in fact a violation of OpenStudy's Code of Conduct. Do you mean log2(3x+20) - log2(x) =3?

OpenStudy (anonymous):

yes

OpenStudy (tyteen4a03):

In that case, remember that logb(m/n) = logb(m) – logb(n).

OpenStudy (anonymous):

Thank you for the formula, I was having a problem with the last part.

OpenStudy (whpalmer4):

Also good to ignore answers without explanations because they may be incorrect :-)

OpenStudy (anonymous):

@tyteen4a03 I'm sorry I wasn't aware of that next time I'll try to explain how. Thanks for letting me know :)